OpenStudy (anonymous):

Which is the formula to calculate area for a circle with diameter 82 miles?

OpenStudy (anonymous):

pi*(41)^2

OpenStudy (btaylor):

radius = 0.5 x diameter = 41 miles area = pi r^2 area = pi (41)^2 area = 1681 pi, about 5281 square miles.